Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
Thanks for any help you can provide for me.

Answer by bam878s(77)

Let x = the number of nickels she has and y = the number of dimes she has. From the information provided, we can see that:
$.05(x) + $.10(y) = $3.50 and
x + y = 50.
From the second of these equations we can solve for x
x + y = 50 ==> x = 50 - y (by subtracting y from both sides)
Now we can plug this value of x back into the first equation above.
.05(50-y) + .10y = 3.50 (multiply the .05 over 50 - y)
[.05(50) - .05(y)] + .10y = 3.50
2.5 - .05y + .10y = 3.50 (add the y terms)
2.5 + .05y = 3.50 (now subtract both sides by 2.5)
.05y = 3.50 - 2.50
.05y = 1 (divide both sides by .05)
y = 20 dimes!
Now we can use the second of the original equations to solve for x!
x + y = 50
x + 20 = 50
x = 50 - 20
x = 30.
So we have 30 nickels!
Lets check!
$0.05(30) + $0.10(20) = $1.50 + $2.00 = $3.50
There are 30 nickels and 20 dimes.
Hope this helps you! Hang in there and don't get discouraged!